OpenStudy (anonymous):

if sin x co(pi/7)-sin(pi/7)cos x= -sqrt2/2 then x can equal ____ check all that apply. a- 3pi/4+pi/7+2npi b- 5pi/4+pi/7+2npi c- pi/4+pi/7+2npi d- 7pi/4+pi/7+2npi

OpenStudy (anonymous):

thats cos not co. typo.

OpenStudy (anonymous):

your this expression becomes Sin(x-pi/7)= -sqrt 1 Sin(x-pi/7)=-1 Sin(x-pi/7)= -sin(pi/2) Sin(x-pi/7)=sin(3pi/4) x= 3pi/4 + pi/7+ 2n pi that is a your answer is a